OpenStudy (anonymous):

Simplify the trigonometric expression. sin^2 theta / 1 - cos theta

OpenStudy (campbell_st):

well use \[\sin^2(\theta) + \cos^2(\theta) = 1\] which can be written as \[\sin^2(\theta) = 1 - \cos^2(\theta)\] so your question becomes \[\frac{1 - \cos^2(\theta)}{1 - \cos(\theta)}\] factor the numerator and you'll see a common factor

OpenStudy (anonymous):

so would the answer be 1+ cos theta

OpenStudy (anonymous):

@campbell_st ??

OpenStudy (freckles):

\[\frac{u^2-v^2}{u-v}=\frac{(u-v)(u+v)}{u-v} \text{ this step factored a difference of squares } \\ =u+v \text{ this step used } \frac{u-v}{u-v}=1 \text{ assumning of course } u \neq v\]

OpenStudy (freckles):

so yes

OpenStudy (campbell_st):

it would be... well done

OpenStudy (anonymous):

thank you !